OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a vertex at the origin and a focus at (0, -4).

OpenStudy (anonymous):

y = -1/4 x^2 y^2 = -4x y^2 = -16x y = -1/16 x^2

hero (hero):

Just a second...

OpenStudy (anonymous):

no problem

hero (hero):

If given two points, the focus \((x_1, y_1)\) and the directrix \((x_2, y_2)\), you can insert them into the following formula: \((x - x_1)^2 + (y - y_1)^2 = (x - x_2)^2 + (y - y_2)^2\) then simplify afterwards to get the standard form of the parabola.

OpenStudy (anonymous):

ok so what am i plugging in?

hero (hero):

Oh, wrong question. You have the vertex and the focus, but not the directrix.

OpenStudy (anonymous):

yea thats it

hero (hero):

There's a process to find the directrix.

OpenStudy (anonymous):

thats where im lost

hero (hero):

Did you know that any point on the parabola is equidistant from the focus and the directrix?

hero (hero):

This means that we can easily find the directrix since we have the vertex.

OpenStudy (anonymous):

so how?

hero (hero):

Well, basically, we know that the focus is (0,-4) and the vertex is (0,0). Since the focus is 4 units below the vertex, that means the directrix is four units above it.

hero (hero):

We also know that the directrix is a horizontal line. In this case it will be y = 4. When we express the directrix as a point, it becomes \((x, 4)\)

OpenStudy (anonymous):

oh ok i understand. so then from there i would plug it in to the equation you gave earlier?

hero (hero):

So now we have the focus (0,-4) and the directrix (x,4). Plug those points into the formula above to find the standard form of the equation of the parabola. Yes.

OpenStudy (anonymous):

do i need to solve for the missing x in the directrix

hero (hero):

You don't. You insert the x into the formula in place of \(x_2\)

OpenStudy (anonymous):

oh ok

hero (hero):

At this point, it's probably a good idea to show the work you've done so far that way I can make sure you've done this step correctly.

OpenStudy (anonymous):

how would i do that? im on the computer and im doing my work on paper...

hero (hero):

Use the draw button or use \(LaTeX\)

OpenStudy (anonymous):

ok ill have to draw it because my computer doesn't support latex

hero (hero):

You can still type the LaTeX. It won't stop it from showing up on my end.

OpenStudy (anonymous):

just give me a couple minutes

hero (hero):

There's a draw button you can click BTW.

OpenStudy (anonymous):

i know im just in the middle of doing the work

hero (hero):